Sam is 6 yrs older than his sister Natalie. The sum of their age is 30 yrs. What is Natalie age?

Answer by checkley79(3341)

S=N+6
S+N=30
(N+6)+N=30
2N=30-6
2N=24
N=24/2
N=12 YEARS OLD FOR NATALIE.
S=12+6=18 YEARS OLD FOR SAM.
PROOF:
12+18=30
30=30